(15 pts) Complete the triangle: Angles A = 290 Sides a- b = 92.1 C= 90° C=
Answer #1

as we know the sum of the the angle of triangle is 180 degree

A+B+C=180

29+b+90=180

119+b=180

b=180-119=61

so angle b is 61 degree

now we can find side a and c by using sine formula

A/sina=B/sinb

A/sin29=92.1/sin61

A=(92.1X0.48481)/0.8746=44.651/0.8746=51.05

SO SIDE A=51.05

now B/sinb=C/sinc

92.1/0.8746=C/sin90

105.305=C/1

SO SIDE C=105.31